1. Reaction of l- with Fe3+ in acidic and basic solution a. Kl(aq) + FeCl3(aq) (ACIDIC) • Observations • Reduction Half React What is the reduction half reaction, oxidation half reaction, overall balanced redox reaction, and the occurence/non-occurebce of reaction?
Answer #1

Answer:

The reaction of I- with Fe3+ in acidic medium

The iodide is oxidised to iodine and Fe3+ is reduced to Fe2+

Thus the reduction half reaction is

      Fe3+ + 1 e-\rightarrow Fe2+   ------------------ (1)

The oxidation half reaction is

    2 I-\rightarrow I2 + 2 e- --------------- (2)

The overall reaction is obtained by multiplying equation (1) by 2 and add to equation (2)

[ This is to balance the number of electrons on both sides.]

2 Fe3+ + 2 I-\rightarrow I2 + 2 Fe2+ ------------------ (3)

This equation (3) is the balanced overall redox reaction

It can be expressed in cell notation as

/oxidation//reduction/            or     /anode cell//cathode cell/

that is 2I- / I2   //   Fe3+ / Fe2+

                               anode    cathode

                             (oxidation)     ( reduction)

E0cell = Eox (anode) + Ered (cathode)

        From the standard reduction potential table

                                        the reduction potential for I2 + 2 e-\rightarrow 2I-       E = 0.54 V

                               Therefore the oxidation potential = -0.54 V

                the reduction potential for Fe3+ + e-\rightarrow Fe2+   E = 0.77 V

Therefore E0cell = -0.54 + 0.77 = + 0.23 V

           

Since E0cell is positive , \DeltaG0 = -nFE0 will be negative and hence the forward reaction is spontaneous.

E0cell for the reverse reaction is -0.23 V

Since E0cell is negative , \DeltaG0 = -nFE0 will be positive and hence the reverse reaction is non-spontaneous.

Thus the reaction

                                 2 Fe3+ + 2 I-\rightarrow I2 + 2 Fe2+ (feasible)

                                  I2 + 2 Fe2+    \rightarrow     2 Fe3+ + 2 I- ( not feasible)

                                 is occurence in the forward direction. The reverse reaction is non-occurence.